# Sparsity-Aware Coding for Single Photon Sensitive Vision Using Selective Sensing

### v4

@@ IV-B Model Validation
Given the universality of reconstruction algorithms, it is imperative that the input data vector $\boldsymbol{x}$, is not compressible. Hence, we generate $\boldsymbol{x}$ by randomly sampling from the uniform distribution $U(0,1)$. Throughout the optimization process, these inputs are regenerated 10 times, thereby introducing additional randomization and reducing compressibility. The output is the reconstructed $\boldsymbol{x}$, with MSE between the input and the output serving as the loss function. The simulations are conducted with a total of $1\times 10^{2}$ photons which is used to rescale the $\boldsymbol{x}$ to simulate a short exposure, given that AGN and Poisson noise behave more differently under low-light conditions. We initialize the masks $\boldsymbol{M}$ with an $N\times N$ identity matrix and a $\{0,1\}$ Hadamard matrix respectively, and optimize them until the model reports constant losses.
Fig. 3(a) displays the first 6 optimized masks from both initial matrices generated by the ONN model under both AGN and Poisson noise. Specifically, it tends to open more pixels to enhance light throughput under AGN, but adheres to the RS mode under the Poisson noise. Fig. 3(b) provides a more detailed visualization of the optimized masks by presenting the distribution of their entries. The x-axis represents the values in the masks, while the y-axis shows the total number of entries within a specific range. This observation aligns with the reconstruction results, where the RS performed modestly under AGN but considerably better under the Poisson noise. Gradients incorporating these two noise models during the training process lead the mask optimization along very different directions. From this example we can see that it is not appropriate to design a coding system under AGN and then apply it to Poisson noise data.
(a) First 6 ONN Optimized Masks
Initial matrix Gaussian Poisson
@@
Initial matrix Gaussian Poisson
@@ IV-C Computer Simulated Experiments
In this section, we present two distinct classification tasks. Initially, we assess the effectiveness of various coding schemes using the MNIST dataset, focusing on their performance in number recognition. Subsequently, we extend the evaluation to hyperspectral data, examining the adaptability and efficacy of the coding schemes in this different domain.
@@ IV-C1 MNIST Simulated Data Preparation
The MNIST dataset, which comprises handwritten digits ranging from 0 to 9, has a default size of 28 by 28 pixels [37]. To align the data with Hadamard matrices, we add black pixels to the edges of the images and resize them to 32 by 32 pixels [37]. Subsequently, we rescale the pixel values from the original range of $[0,1]$ to $[0.3,1]$. This rescaling introduces non-zero Poisson noise to the black regions, which has variance proportional to the expected photon counts.
We evaluate our proposed model through simulations incorporating both AGN and Poisson noise models, using varying exposure times to generate results at different noise levels. We also test all compressible strategies, varying the total number of masks but fixing the total exposure time, and define a metric called compression ratio as the ratio of the number of masks to the number of pixels [38]. For RS and HB, this ratio is constrained to be 1.00, while the rest are evaluated over values in $\mathcal{C}_{R}=\{0.01,0.04,0.09,0.16,0.25,0.36,0.49,0.64,0.81,1.00\}$ so that $\boldsymbol{M}$ of each coding scheme is not restricted to be full-rank. The final comparison only involves the best results among all compression ratios for each strategy. Unlike pre-defined masks used by other strategies, the initial masks for ONN can affect performance in the presence of noise. We therefore initialized its masks with the PCA components from the training data.
@@ IV-C2 Spectral Datasets and Performance Evaluation
We utilize the Indian Pines dataset [39], acquired through AVIRIS on June 12, 1992, covering Purdue University’s Agronomy farm and surrounding areas. With $145\times 145$ pixels and 224 spectral bands (0.4 to 2.5 $\mu m$) per pixel, this dataset offers a comprehensive view of this area. Fig. 4(a) illustrates 16 classes in Indian Pines, along with their counts and semantic labeling ground truth. It features a representation of the band-12 for all pixels. Fig. 4(b) shows the relative intensity of the selected classes from the spectrum data, providing a concise overview of their differences.
@@
In the experiment, spectral data vectors, extracted as arrays from each pixel, are zero-padded from 224 to 256 in length for Hadamard mask alignment and rescaled to $[0,1]$. It undergoes processing using the modified scanner-classifier module illustrated in Fig. 2. The classifier is streamlined to a single fully-connected layer with 1478 neurons, utilizing the default architecture provided by scikit-learn during the pre-project testing phase which gives an 86% classification rate without any coding or noise. The ONN classification layers are adjusted to accommodate input data subsets, each containing 400 entries per class. Classes with less than 400 entries are omitted to maintain unbiased input.
Model performance is evaluated using average classification rates from five independent tests. Each test involves a random 90%-10% training-validation split of the dataset. Training continues until the validation rate plateaued or declined, with the peak validation rate representing the upper bound for each test. The final performance metric is the average of all five peak validation rates, ensuring a reliable and accurate assessment of the model.
@@ IV-D Hardware Experiments
To further support our conclusions, we conduct an experiment in addition to simulations. As shown in Fig. 5, the experiment utilizes a DMD (digital micromirror device), specifically the DLP 7000 model from Texas Instruments, and a PMT (photomultiplier tube) from PicoQuant, specifically the PMA Series. The experimental setup involves both hardware-end data acquisition and software-end classifier training. Handwritten images are displayed on a monitor, captured by this vision system, and subsequently identified according to the depicted numbers.
@@
To ensure that the classifier is properly trained using the PMT data, we perform a transformation $g(\cdot)$ incorporating illumination correction and calibration of the MNIST images $\boldsymbol{X}$ to match what the PMT would actually ”observe”, $\boldsymbol{X}^{g}$, during data acquisition. Mathematically, this transformation can be described as $\boldsymbol{X}^{g}=g(\boldsymbol{X})$. This step aims at reducing the model’s inconsistency between between optimizing the sensing matrices $\boldsymbol{M}$ and classifiers with digital data on a computer and testing the models with acquired data from experiments. Specifically, we perform a raster scan of a sample from $\boldsymbol{X}_{s}$, a random subset from $\boldsymbol{X}$ , with each mask exposed for 1 second. The light intensity is adjusted to ensure that the brightest pixel generated approximately 1000 photon counts per second. The data $\boldsymbol{X}_{s}^{g}$ captured during this experiment is then reshaped into a 32 by 32 image, as shown in Fig. 6, and $g(\cdot)$ is then determined. Applying the transformation to digital data yields $\boldsymbol{X}^{g}$, which is then fed into models featuring various coding schemes for performance evaluation.
Following the generation of mask pattern sequences, they are uploaded onto the DMD to capture data, represented as $\tilde{\boldsymbol{y}}$. Our test set consists of ten handwritten images, each of which appears exactly once during the experiment. To control the noise level in the data, we varied lengths of random intervals throughout the 1-second exposure time for each mask. For instance, by considering only the photons detected in the first 0.5-second interval, we effectively double the noise level. By preserving the arrival times of all photons, we are able to systematically adjust the level of noise in the acquired data.
@@ V-B Classification Performances on Hardware Experiment Data
In our experiments, the collected data takes the form of a series of timestamps recording the photons’ arrival time. During data collection, both RS and HB uses 1024 masks, while TH and HB employ the same measured data, but TH uses only the first 92 measurements for compression purposes. Similarly, PCA and ONN utilizes 92 masks. Physical measurements take one second per mask for all strategies. To generate data with higher noise levels from the raw data, we randomly select intervals within the 1-second span and count the total number of photons within them. Since ONN and PCA have only 92 masks, the maximum exposure time is set to 92 seconds. To ensure a fair comparison, we compute the interval length for each strategy as $\frac{\tau}{m}$, where $\tau$ represents the total exposure time under investigation, and $m$ denotes the number of masks used for each strategy.
@@
Fig. 8 displays the classification rates achieved on the experimental data. The x-axis shows the exposure time, and the y-axis represents the classification rates of the models. Although there may be discrepancies between software training and hardware data acquisition, the ONN models outperform other coding schemes across all noise levels. However, the figure reveals some unexpected phenomena. Firstly, the HB strategy outperforms the RS for short exposure times, which is not observed in the simulations. This phenomenon may have arisen due to the dark counts in the PMT, a signal-independent noise that has about 40-80 counts per second under our lab condition. Secondly, TH achieves better classification performances than PCA when the exposure time is short. This result is due to the use of unneeded PCA masks since the compression ratio was not optimized as it was in the simulation. To verify, we conduct simulations and obtain Fig. 9 by fixing the total number of photons at $10^{5}$ and varying the compression ratio under the Poisson noise. This figure shows that PCA method can perform worse than the TH if an improper compression ratio is chosen. Hence, PCA strategy requires careful estimation of the noise level to avoid improper compression ratios. Notably, ONN demonstrates enhanced stability and robustness when confronted with this challenge. This flexibility in application allows for broader use of the ONN beyond its specific training conditions, increasing its potential impact and utility in real-world scenarios.
@@
Figure 9: Classification rates on simulated data w.r.t compression ratios and light levels. The models are trained under Poisson noise using a light level of $10^{5}$ photons and subsequently tested under different light levels. It is observed that hardware PCA performance deteriorates when the compression ratio is not properly chosen, whereas the Truncated Hadamard and ONN models exhibit greater robustness in such scenarios.
@@ VI Discussion and Limitations
Despite the promising results of our project, there are some limitations that must be acknowledged. First, we used a Gaussian noise model with reparameterization to train our model, which is only an approximation of the actual Poisson noise at the sensor. Second, our test set consisted of only 10 numbers, which may not provide a comprehensive evaluation of the model’s performance. These limitations highlight the need for further improvements in the optimization of the ONN model, such as using more advanced optimization methods and larger sets of experimental data. Our work highlights the importance of the integration of imaging hardware and signal processing. In single photon accurate imaging systems, comprehensibility and sparsity of the data can be exploited to far greater effect during the measurement, as opposed to post processing.
@@ Acknowledgment
The authors would like to thank Dr. Rebecca Willet for her invaluable assistance in refining the problem statement and providing support for our conclusions during the course of this research. Additionally, the authors would like to express gratitude to Dr. Sebastian Bauer for his guidance in understanding the mathematical aspects of this problem at its early stages. This work was supported by funding from the National Science Foundation Career grant.

### v8

@@ IV-B Model Validation
Given the universality of reconstruction algorithms, it is imperative that the input data vector $\boldsymbol{x}$, is not compressible. Hence, we generate $\boldsymbol{x}$ by randomly sampling from the uniform distribution $U(0,1)$. Throughout the optimization process, these inputs are regenerated 10 times, thereby introducing additional randomization and reducing compressibility. The output is the reconstructed $\boldsymbol{x}$, with MSE between the input and the output serving as the loss function. The simulations are conducted with a total of $1\times 10^{2}$ photons which is used to rescale the $\boldsymbol{x}$ to simulate a short exposure, given that AGN and Poisson noise behave more differently under low-light conditions. We initialize the masks $\boldsymbol{M}$ with an $N\times N$ identity matrix and a $\{0,1\}$ Hadamard matrix, respectively, and optimize them until the model reports constant losses.
Fig. 3(a) displays the first 6 optimized masks from both initial matrices generated by the ONN model under both AGN and Poisson noise. Specifically, it tends to open more pixels to enhance light throughput under AGN, but adheres to the RS mode under Poisson noise. Fig. 3(b) provides a more detailed visualization of the optimized masks by presenting the distribution of their entries. The x-axis represents the values in the masks, while the y-axis shows the total number of entries within a specific range. This observation aligns with the reconstruction results, where the RS performed modestly under AGN but considerably better under Poisson noise. Gradients incorporating these two noise models during the training process lead mask optimization in very different directions. From this example, we can see that it is not appropriate to design a coding system under AGN and then apply it to Poisson noise data.
(a) 6 Random Selected ONN Optimized Masks
Gaussian Poisson
@@
Gaussian Poisson
@@ IV-C Computer Simulated Experiments
In this section, we present two distinct classification tasks. Initially, we evaluated the effectiveness of various coding schemes using the MNIST dataset, focusing on their performance in number recognition. Subsequently, we extend the evaluation to hyperspectral data, examining the adaptability and efficacy of the coding schemes in this different domain.
@@ IV-C1 MNIST Simulated Data Preparation
The MNIST dataset, which consists of handwritten digits ranging from 0 to 9, has a default size of 28 by 28 pixels [37]. To align the data with Hadamard matrices, we add black pixels to the edges of the images and resize them to 32 by 32 pixels [37]. Subsequently, we rescale the pixel values from the original range of $[0,1]$ to $[0.3,1]$. This rescaling introduces non-zero Poisson noise to the black regions, which has variance proportional to the expected photon counts.
We evaluate our proposed model through simulations incorporating both AGN and Poisson noise models, using varying exposure times to generate results at different noise levels. We also test all compressible strategies, varying the total number of masks but fixing the total exposure time, and define a metric called compression ratio as the ratio of the number of masks to the number of pixels [38]. For RS and HB, this ratio is restricted to 1.00, while the rest are evaluated over values in $\mathcal{C}_{R}\in\{0.01,0.04,0.09,0.16,0.25,0.36,0.49,0.64,0.81,1.00\}$ so that $\boldsymbol{M}$ of each coding scheme is not restricted to being full-rank. The final comparison involves only the best results among all compression ratios for each strategy. Unlike pre-defined masks used by other strategies, the initial masks for ONN can affect performance in the presence of noise. We therefore initialized its masks with the PCA components from the training data.
@@ IV-C2 Spectral Datasets and Performance Evaluation
We use the Indian Pines dataset [39], acquired through AVIRIS on June 12, 1992, covering the Purdue University Agronomy farm and surrounding areas. With $145\times 145$ pixels and 224 spectral bands (0.4 to 2.5 $\mu m$) per pixel, this dataset offers a comprehensive view of this area. Fig. 4(a) illustrates 16 classes in Indian Pines, along with their counts and semantic labeling of the ground truth. It features a representation of band-12 for all pixels. Fig. 4(b) shows the relative intensity of the selected classes from the spectrum data, providing a concise overview of their differences.
@@
In the experiment, spectral data vectors, extracted as arrays from each pixel, are zero-padded from 224 to 256 in length for Hadamard mask alignment and rescaled to $[0,1]$. It undergoes processing using the modified scanner-classifier module illustrated in Fig. 2. The classifier is simplified into a single fully connected layer with 1478 neurons, using the default architecture provided by scikit-learn during the preproject testing phase which gives an 86% classification rate without any coding or noise. The ONN classification layers are adjusted to accommodate input data subsets, each containing 400 entries per class. Classes with fewer than 400 entries are omitted to maintain an unbiased input.
Model performance is evaluated using average classification rates from five independent tests. Each test involves a random 90%-10% training-validation split of the dataset. Training continues until the validation rate plateaued or declined, with the highest validation rate representing the upper bound for each test. The final performance metric is the average of the five maximum validation rates, ensuring a reliable and accurate assessment of the model.
@@ IV-D Hardware Experiments
To further support our conclusions, we conduct an experiment in addition to simulations. As shown in Fig. 5, the experiment utilizes a DMD (digital micromirror device), specifically the DLP 7000 model from Texas Instruments, and a PMT (photomultiplier tube) from PicoQuant, specifically the PMA Series. The experimental setup involves both hardware-end data acquisition and software-end classifier training. Handwritten number images are displayed on a monitor, captured by this vision system, and subsequently identified according to the depicted numbers.
@@
To ensure that the classifier is properly trained using the PMT data, we perform a transformation $g(\cdot)$ that incorporates illumination correction and calibration of the MNIST images $\boldsymbol{X}$ to match what the PMT would actually ”observe”, $\boldsymbol{X}^{g}$, during data acquisition. Mathematically, this transformation can be described as $\boldsymbol{X}^{g}=g(\boldsymbol{X})$. This step aims at reducing the model inconsistency between optimizing the sensing matrices $\boldsymbol{M}$ and classifiers with digital data on a computer and testing the models with the data acquired from the experiments. Specifically, we perform a raster scan of a sample from $\boldsymbol{X}_{s}$, a random subset from $\boldsymbol{X}$ , with each mask exposed for 1 second. The light intensity is adjusted to ensure that the brightest pixel generated approximately 1000 photon counts per second. The data $\boldsymbol{X}_{s}^{g}$ captured during this experiment is then reshaped into a 32 by 32 image, as shown in Fig. 6, and $g(\cdot)$ is then determined. Applying the transformation to digital data yields $\boldsymbol{X}^{g}$, which is then fed into models featuring various coding schemes for performance evaluation.
Following the generation of mask pattern sequences, they are uploaded onto the DMD to capture data, represented as $\tilde{\boldsymbol{y}}$. Our test set consists of ten handwritten images, each of which appears exactly once during the experiment. To control the noise level in the data, we varied lengths of random intervals throughout the 1-second exposure time for each mask. For instance, considering only the photons detected in the first 0.5-second interval, we effectively double the noise level. By preserving the arrival times of all photons, we can systematically adjust the level of noise in the acquired data.
@@ V-B Classification Performances on Hardware Experiment Data
In our experiments, the collected data take the form of a series of timestamps recording the photon arrival time. During data collection, both RS and HB uses 1024 masks, while TH and HB employ the same measured data, but TH uses only the first 92 measurements for compression purposes. Similarly, PCA and ONN utilizes 92 masks. Physical measurements take one second per mask for all strategies. To generate data with higher noise levels from the raw data, we randomly select intervals within the 1-second span and count the total number of photons within them. Since ONN and PCA have only 92 masks, the maximum exposure time is set to 92 seconds. To ensure a fair comparison, we compute the interval length for each strategy as $\frac{\tau}{m}$, where $\tau$ represents the total exposure time under investigation, and $m$ denotes the number of masks used for each strategy.
@@
Fig. 8 displays the classification rates achieved on the experimental data. The x-axis shows the exposure time, and the y-axis represents the classification rates of the models. Although there may be discrepancies between software training and hardware data acquisition, the ONN models outperform other coding schemes across all noise levels. However, the figure reveals some unexpected phenomena. Firstly, the HB strategy outperforms the RS for short exposure times, which is not observed in the simulations. This phenomenon may have arisen due to the dark counts in the PMT, a signal-independent noise that has about 40-80 counts per second under our lab condition. Secondly, TH achieves better classification performances than PCA when the exposure time is short. This result is due to the use of unneeded PCA masks since the compression ratio was not optimized as it was in the simulation. To verify, we conduct simulations and obtain Fig. 9 by fixing the total number of photons at $10^{5}$ and varying the compression ratio under Poisson noise. This figure shows that PCA method can perform worse than the TH if an improper compression ratio is chosen. Hence, PCA strategy requires careful estimation of the noise level to avoid improper compression ratios. Notably, ONN demonstrates enhanced stability and robustness when confronted with this challenge. This flexibility in application allows for broader use of the ONN beyond its specific training conditions, increasing its potential impact and utility in real-world scenarios.
@@
Figure 9: Classification rates on simulated data regarding compression ratios and light levels. The models are trained under Poisson noise using a light level of $10^{5}$ photons and subsequently tested under different light levels. It is observed that hardware PCA performance deteriorates when the compression ratio is not properly chosen, whereas the Truncated Hadamard and ONN models exhibit greater robustness in such scenarios.
@@ VI Discussion and Limitations
Despite the promising results of our project, there are some limitations that must be acknowledged. First, we used a noise model where the variance equals the mean and applied reparameterization to approximate the actual photon noise at the sensor. This approach offers a valid alternative to Poisson noise, which is also just an approximation. Second, our test set consisted of only 10 numbers, which may not provide a comprehensive evaluation of the model’s performance. These limitations highlight the need for further improvements in the optimization of the ONN model, such as using more advanced optimization methods and larger sets of experimental data. Our work highlights the importance of the integration of imaging hardware and signal processing. In single photon accurate imaging systems, comprehensibility and sparsity of the data can be exploited to far greater effect during the measurement, as opposed to post processing.
@@ Acknowledgment
The authors would like to thank Dr. Rebecca Willet for her invaluable assistance in refining the problem statement and providing support for our conclusions during the course of this research. Additionally, the authors would like to express gratitude to Dr. Sebastian Bauer for his guidance in understanding the mathematical aspects of this problem at its early stages. This work is supported by the National Science Foundation (1846884) and the Air Force Office for Scientific Research (FA9550-21-1-0341).